OpenStudy (anonymous):

a copy machine makes 24 copies per minute how many copies does it make in 4 minutes and 15 seconds!!!:)

OpenStudy (anonymous):

The easiest way to do this is cross multiply so you do 24/1=x/4.15

OpenStudy (lizzyvet1):

102 copies 24/4=6 24 x 4 =96 96 + 6 =102

OpenStudy (anonymous):

let x- be copies copied in one minute x=24 since there are 60 seconds in one minute 4+15/60 =17/4 17/4 (x) 17/4 (24) =408/4 =102 copies